OpenStudy (anonymous):

Simplify the product using the rules for exponents: (142)6

OpenStudy (anonymous):

(14^2)^6

OpenStudy (anonymous):

\[\huge (a^b)^c = a^{b \times c}\]

OpenStudy (shane_b):

Since you're raising an exponent to an exponent, you keep the same base and multiply the exponents. The rule is:\[\large (a^x)^y=a^{xy}\]

OpenStudy (anonymous):

So in this case b is 2 and c is 6..